# SyncDiff: Synchronized Motion Diffusion for Multi-Body Human-Object Interaction Synthesis

### v2

@@ B.1 Hyperparameters for Frequency-based Decomposition and Explicit Synchronization
[…]
We conduct experiments on val1 split from TACO using different $t_{0}$, as shown in Table  6. RA refers to recognition accuracy. In practice, we choose $t_{0}=50$.
$t_{0}$ Inference time per sample(s) RA(%, $\uparrow$)
@@
Table 6: Results for different $t_{0}$ on val1 split from TACO.
B.2 A Brief Introduction to BPS Algorithm
B.3 Hyperparameters in Model Architecture
Action/object Label Feature Extraction Branch. Specific parameters are shown in Table  7.
Component Description
CLIP Type ViT-B/32
Raw Feature Space Dimension $512$
MLP Architecture Linear($512$, $512$)
ReLU()
Linear($512$, $128$)
Table 7: Hyperparameters in label feature extraction branch.
Object Geometry Feature Extraction Branch. Specific parameters are shown in Table  8.
Component Description
Raw Feature Space Dimension $1024\times 3$
MLP Architecture Linear($1024\times 3$, $512$)
ReLU()
Linear($512$, $128$)
Table 8: Hyperparameters in rigid body shape feature extraction branch.
Transformer Encoder-Decoder. After extracting all label features and the geometric features of rigid bodies, we concatenate them with the shape parameters of articulated skeletons to form a condition vector, whose shape is $\mathbb{R}^{128\times(2m+1)+10n}$, as is mentioned in Section  4.2.1. After replicating and concatenating with $\bar{x}$ and $\text{AC}(x)$, the dimension becomes $\mathbb{R}^{T\times\left(128\texttimes(2m+1)+10n+7m+3Dn+7{m\choose 2}+3Dmn\right)}=\mathbb{R}^{T\times\left(128\texttimes(2m+1)+10n+D_{\text{sum}}\right)}=\mathbb{R}^{T\times\mathcal{C}}$. The architecture of the transformer encoder-decoder is shown in Table  9.
Component Description
Encoder Conv1D($\mathcal{C}$, $512$)
Latent Transformer $4$-layer, $8$-head, $1024$-dim
Decoder Linear($512$, $D_{\text{sum}}$)
Table 9: Hyperparameters in the transformer encoder-decoder.
B.4 Training and Inference Hyperparameters
Other important hyperparameters for training and inference process are shown in Table  10.
Parameter TACO CORE4D OAKINK2 GRAB
Batch Size $32$
[…]
Epoch $250$k $140$k $280$k $100$k
$\lambda_{\text{DC}}$ $1$ $1$ $1$ $1$
$\lambda_{\text{AC}}$ $0.8$ $0.3$ $0.3$ $0.1$
$\lambda_{\text{norm}}$ $0.1$ $0.1$ $0.1$ $0.1$
$\lambda_{\text{align}}$ $0.3$ $0.3$ $0.3$ $0.15$
$\lambda_{\text{exp}}$ $0.3$ $0.3$ $0.3$ $0.3$
Diffusion $\alpha\in[0.0001,0.01]$, Uniform
B.5 Hyperparameters for Mesh Reconstruction
Hand Mesh Reconstruction. Specific parameters for the three hand-object interaction datasets are shown in Table  11.
@@ B.5 Hyperparameters for Mesh Reconstruction
Table 11: Hyperparameters for hand mesh reconstruction.
Hand Mesh Reconstruction. Specific parameters for the human-object interaction dataset CORE4D are shown in Table  12.
Parameter CORE4D
@@
Epoch $5$k
$\lambda_{\text{pos}}$ $3$
$\lambda_{\text{vel}}$ $0.1$
Table 12: Hyperparameters for human body mesh reconstruction.
Appendix C Time and Space Cost, Hardware Configurations
Time and Space Cost for Training and Inference. The training time, average inference time per sample, and GPU memory usage during training are detailed in Table  13.
Cost TACO CORE4D OAKINK2 GRAB
Training Time $20.7$h $9.5$h $40.1$h $7.9$h
Inference Time $7.7$s $6.5$s $7.3$s $6.6$s
Memory Cost $6.11$G $5.07$G $8.59$G $4.93$G
Time and Space Cost for Mesh Reconstruction. Time and space cost of performing mesh reconstruction for a motion sequence of $T=200$ frames are shown in Table  14.
Cost TACO CORE4D OAKINK2 GRAB
Time Cost $130$s $144$s $206$s $130$s
Memory Cost $492$M $826$M $492$M $492$M
Appendix D Experimental Details
D.1 Dataset Statistics
The sizes of each dataset split are shown in Table  15.
Dataset Statistics
TACO train:test1:test2:test3:test4=1035:238:260:403:610
CORE4D train:test1:test2=483:197:195
OAKINK2 train:val:test=1884:167:723
GRAB train:val:test=992:198:144(Unseen Subject)
train:test=1126:208(Unseen Object)
Table 15: Dataset statistics.
D.2 Pseudocode for CSR, CRR, and CSIoU
In this section, we will provide a detailed description of three physics-based metrics: CSR, CSIoU, and CRR.
First, we define two types of contact: surface contact and root contact. Here $o$ is one single object mesh sequence of size $\mathbb{R}^{T\times M\times 3}$, where $M$ is the number of vertices on its mesh. $h$ is hand mesh sequence of size $\mathbb{R}^{T\times 778\times 3}$ in Contact_Surface, while it denotes trajectories of root joints of two hands of size $R^{T\times 2\times 3}$ in Contact_Root, as shown in Algorithm 2.
@@ Appendix E Limitations and Discussions
Some limitations of SyncDiff and their potential solutions are as follows:
@@
1. Lack of Articulation-Aware Modeling. Our method models articulated objects as part-wise rigid body individuals directly and coordinates their motions without leveraging their intrinsic articulations. Integrating these articulations into multi-body likelihood modeling could be an interesting future direction.
@@
2. High-cost of Explicit Synchronization Step. Our explicit synchronization step still incurs a certain time consumption. As the number of bodies increases, not all pairwise relationships are necessary. A possible solution is to use another learning-based method to filter out the relationships that truly require synchronization.
@@
3. Lack of Physically Accurate Guarantees. Unlike methods that utilize true physical simulations, our approach cannot guarantee physical truthfulness. In many cases, minor errors can be observed in the supplementary videos, but these small discrepancies may be sufficient to cause visible failures in real tasks. We prefer to treat SyncDiff as a robot planning method that requires downstream integration with physically accurate optimization to ensure practical usability in real-world applications.
4. Unstable Grasp for Objects with Complex Structures. For objects like chairs and desks, which have many different functional parts, BPS occasionally fails to effectively extract surface information for the objects. This results in unstable grasping, especially when the geometric features of the objects are not present in the training set, as shown in Figure  8. This might be solved by further decomposing rigid objects into rigid functional parts with relatively simple topologies.
Figure 8: Qualitative results from CORE4D [99] dataset.
5. Irrelevant Body Trajectory Interference. For irrelevant bodies that have already completed the interaction or have not yet joined the interactions, SyncDiff sometimes fails to effectively plan a joining or leaving way for them, leading to interference with subsequent interactions. We provide three samples from CORE4D in the Failure Cases section in the supplementary visualization results for illustration.
6. Rotation Instability. $SO(3)$ is not isomorphic to Euclidean space, and quaternions sometimes fail to maintain stability, especially after introducing decomposition based on frequencies. This usually occurs in cases where objects need to undergo large rotations, such as some types of tools in TACO. We provide three samples from TACO in the Failure Cases section in the supplementary visualization results for illustration.
7. Limitations in Pure Multi-human Interaction Synthesis. Since our relative representations need to be generated in the rigid body coordinate system, our method may not be adapted for pure multi-human interaction synthesis. However, by designing more complex relative representations, this limitation could potentially be addressed.

### v3

@@ B.1 Determine the cutoff boundary $L$ for Frequency Decomposition
Since $L=16$ is already enough to almost perfectly reconstruct the motion sequence, to strike a balance between motion representation fidelity and filtering of too high-frequency noise in mocap datasets, in practice, we take $L=16$.
B.2 Determine the Interval $s$ for Explicit Synchronization
In explicit synchronization (Section  3.6), to balance efficiency and performance, we choose a hyperparameter $s(s<<T)$, where $T=1000$ is the total number of denoising steps, which means we only perform explicit synchronization operations every $s$ step, $R=T/s$ times in total.
We conduct experiments on TACO Split 1 using different $s$, as shown in Table  7. RA refers to recognition accuracy. To reduce computational cost without affecting the semantic accuracy of synthesized motions too much, in practice, we choose $s=50$.
$s$ Inference time per sample(s) RA(%, $\uparrow$)
@@ B.2 Determine the Interval $s$ for Explicit Synchronization
Table 7: Results for different $s$ on TACO Split 1.
B.3 Check the Reconstruction Quality of High-frequency Components
From the comparison between Ours and “w/o decompose” in the main text and supplementary videos, we can clearly see that if high-frequency components are not explicitly modeled, the motion trajectory tends to miss subtle high-frequency details, which are crucial for the accuracy of motion semantics. For example, the spatula merely contacts and gets stuck on the plate instead of moving back and forth to complete the scrape off motion; similarly, when a person walks, his/her legs fail to alternate properly and instead slide forward like “ice skating”.
However, frequency decomposition consists of 1) the filter of too high-frequency noise, and 2) the explicit frequency domain representation ($x_{\text{F}}$) of high-frequency components with semantics ($x_{\text{ac}}$), where the former is relatively trivial. We need to decouple the contributions of them in further ablation studies. Additionally, a noteworthy phenomenon is that, compared to the baselines, our method exhibits less jitter (See videos of CORE4D/OAKINK2/GRAB “comparison to baselines”), which could be attributed to either the filter of too high-frequency noise or the two synchronization mechanisms. It is essential to investigate the core source of this effect.
Identifying the source of less jitter is relatively straightforward. It is noted that the motion trajectories for training in the “w/o decompose” scenario includes those too high-frequency noise (See Section  C.4). However, as observed in the “w/o decompose” videos of TACO, there is almost no large-scale jitter in the synthesized motions, with the only issue being the absence of high-frequency interactions between objects. This is because, in the “w/o decompose” experiments, both synchronization mechanisms are effective, ensuring that the generated motions are sufficiently aligned and thus reducing jitter. Conversely, in the OAKINK2 ablation studies of synchronization, whether it’s “w/o align loss” or “w/o exp sync,” the use of frequency decomposition also removes too high-frequency noise, yet it still results in greater jitter compared to the complete SyncDiff (Ours). This sufficiently demonstrates that the contribution to less jitter is primarily due to the synchronization mechanisms, rather than the filter of too high-frequency noise. A more straightforward method of verification is to observe that the errors in Table  6 caused by the filter of too high-frequency noise are almost negligible compared to the scales of the motions, indicating that this operation theoretically has minimal impact on the model’s performance.
To decouple the effects of the frequency-domain based representation and the filter of too high-frequency noise from a more strict aspect, we conduct the following experiments. Besides “w/o decompose”, we add “simply filter”, which replaces $x_{t}$ in “w/o decompose” with $x_{t,\text{dc}}+x_{t,\text{ac}}$, simulating the effect of filtering too high-frequency noise. “Only dc” replaces $x_{t}$ with $x_{t,\text{dc}}$, simulating the effect of using no high-frequency signals (including those with semantics).
Experiment 1: On the four test splits of TACO, we only add noise for no larger than $200$ steps (where the full train pipeline needs $1000$ steps), and examine the reconstruction ability of different methods. For any predicted motion trajectory $\hat{x}$, we decompose it as described in Section  3.3, to obtain the high-frequency component $\hat{x}_{\text{ac}}$. We then examine the discrepancy between $\hat{x}_{\text{ac}}$ and the ground truth $x_{\text{ac}}$. The experimental results on the four splits of TACO are shown in Table  8, proving that explicitly representing high-frequency signals with semantics in the frequency domain indeed improves the effect of reconstruction.
Error (cm, $\downarrow$)
Method Test1 Test2 Test3 Test4
MACS [73] 2.43 2.37 3.11 2.74
DiffH2[13] 1.74 2.08 3.12 3.31
Ours 1.99 1.63 2.24 2.25
w/o decomp 2.68 3.28 4.12 3.72
simply filter 3.23 3.28 3.64 3.77
only dc 5.42 4.87 6.17 5.89
Table 8: Reconstruction of high-frequency components on TACO [48] dataset. The best in each column is highlighted in bold.
Experiment 2: We test the semantic quality of synthesized motions on the four test splits of TACO. The experimental setting is the same as Table  1 in the main text. Results are shown in Table  9. We can clearly see that frequency decomposition enhances the semantic quality of the synthesized motion, and this can not be achieved by merely removing too high-frequency noise.
FID ($\downarrow$) RA (%, $\downarrow$)
Method Test1 Test2 Test3 Test4 Test1 Test2 Test3 Test4
MACS [73] 10.56 23.24 32.18 42.37 58.40 53.08 33.00 19.02
DiffH2[13] 4.34 17.04 24.92 39.20 61.40 56.70 43.67 28.15
Ours 2.70 2.68 22.96 30.23 73.28 85.92 46.90 40.12
w/o decomp 6.44 21.21 28.67 49.58 56.60 51.85 40.02 22.18
simply filter 7.87 20.38 29.35 37.69 54.71 53.08 39.66 22.34
only dc 42.83 54.62 88.47 85.02 30.26 31.44 25.86 19.74
Table 9: Semantic quality of ablation studies on TACO [48] dataset. The best in each column is highlighted in bold.
Appendix C Eliminated Details in the Main Text
C.1 Formulas for rel and comb
In Section  3.5,  3.6, for two distinct bodies a and b, we define $\text{rel}(x_{\textbf{a}},x_{\textbf{b}})$ as b’s motion relative to a, and let $\text{comb}(x_{\textbf{a}},x_{\textbf{b}\rightarrow\textbf{a}})$ utilize the individual motion of a and relative motion between b and a to compute b’s motion. The detailed expressions are shown below:
If $\textbf{a}=o_{j_{1}}$ and $\textbf{b}=o_{j_{2}}$ are two rigid objects, whose motions are $x_{o_{j_{1}}}=[\textbf{t}_{o_{j_{1}}},\textbf{q}_{o_{j_{1}}}]$ and $x_{o_{j_{2}}}=[\textbf{t}_{o_{j_{2}}},\textbf{q}_{o_{j_{2}}}]$, denote $x_{o_{j_{2}}\rightarrow o_{j_{1}}}=[\textbf{t}_{o_{j_{2}}\rightarrow o_{j_{1}}},\textbf{q}_{o_{j_{2}}\rightarrow o_{j_{1}}}]$, then
$\text{rel}(x_{o_{j_{1}}},x_{o_{j_{2}}})=\left[\textbf{q}_{o_{j_{1}}}^{-1}(\textbf{t}_{o_{j_{2}}}-\textbf{t}_{o_{j_{1}}}),\textbf{q}_{o_{j_{1}}}^{-1}\textbf{q}_{o_{j_{2}}}\right],$ (18)
$\text{comb}(x_{o_{j_{1}}},x_{o_{j_{2}}\rightarrow o_{j_{1}}})=[\textbf{q}_{o_{j_{1}}}\textbf{t}_{o_{j_{2}}\rightarrow o_{j_{1}}}+\textbf{t}_{o_{j_{1}}},\textbf{q}_{o_{j_{1}}}\textbf{q}_{o_{j_{2}}\rightarrow o_{j_{1}}}].$ (19)
If $\textbf{a}=o_{j}$ is rigid object, and $\textbf{b}=h_{i}$ is an articulated skeleton, let $x_{o_{j}}=[\textbf{t}_{j},\textbf{q}_{j}]$, and denote the motion for one of the joints in $x_{h_{i}}$ as $\textbf{p}_{h_{i}}$, the motion for one of the joints in $x_{h_{i}\rightarrow o_{j}}$ as $\textbf{p}_{h_{i}\rightarrow o_{j}}$. We have
$\text{rel}(x_{o_{j}},\textbf{p}_{h_{i}})=\textbf{q}_{j_{1}}^{-1}(\textbf{p}_{h_{i}}-\textbf{t}_{j_{1}}),$ (20)
$\text{comb}(x_{o_{j}},\textbf{p}_{h_{i}\rightarrow o_{j}})=\textbf{q}_{o_{j}}\textbf{p}_{h_{i}\rightarrow o_{j}}+\textbf{t}_{o_{j}}.$ (21)
Note that a must be a rigid object, in order to define a coordinate system based on its transformation matrix. More details can be found in the second question of Section  E.1. Also, $\text{rel}\left(\hat{x}_{\textbf{a}},\hat{x}_{\textbf{b}}\right)$ is different from $\hat{x}_{\textbf{b}\rightarrow\textbf{a}}$. The former is calculated based on the predicted motion of a and b, which should approach $\hat{x}_{\textbf{b}\rightarrow\textbf{a}}$ in order to foster synchronization. The latter is directly predicted by the diffusion model, which should fit the distribution of ground-truth $x_{\textbf{b}\rightarrow\textbf{a}}$, in order to increase data fidelity. Similar arguments hold for comb.
C.2 Post Process: Mesh Reconstruction
For tasks such as robot learning, the joint information for articulated skeletons (hands/humans) is already sufficient. Post processing methods like Reinforcement Learning (RL) or Imitation Learning (IL) can further guide the model to generate physically realistic data with the help of physics-based simulations. Therefore, in such scenarios, we can direct use $\hat{x}_{h_{i\in[1,n]}}$ as predicted results. However, for tasks like animation production or VR/AR, it is necessary to reconstruct the full meshes. In this section, we introduce how to reconstruct MANO hand mesh or SMPL-X human body mesh of natural shape from the predicted joint positions $\mathbb{R}^{N\times D\times 3}$, where $N$ and $D$ are the number of frames and joints.
MANO Hand Mesh Reconstruction. For one single hand, suppose the predicted joint positions are $\hat{x}\in\mathbb{R}^{N\times 21\times 3}$. We start with tunable MANO [68] parameters for joint pose $\theta\in\mathbb{R}^{N\times 45}$, global orientation $R\in\mathbb{R}^{N\times 3}$ and translation $l\in\mathbb{R}^{N\times 3}$, which are all set to zero tensors initially. We freeze hand shape $\beta\in\mathbb{R}^{10}$. During the optimizing process, let the current calculated joint positions be $K\in\mathbb{R}^{N\times 21\times 3}$ based on $\theta$, $R$, and $l$. We want to minimize
Here
$\mathcal{L}_{\text{pos}}=\sum\limits_{t=1}^{N}\sum\limits_{i=1}^{21}\|\hat{x}_{t,i}-K_{t,i}\|_{2}^{2}$
$\mathcal{L}_{\text{angle}}=\sum\limits_{t=1}^{N}\sum\limits_{i=1}^{45}[\max(\theta_{t,i}-u_{i},0)+\max(d_{i}-\theta_{t,i},0)]$
$\mathcal{L}_{\text{vel}}=\sum\limits_{t=1}^{N-1}\|l_{t}-l_{t+1}\|_{2}^{2}$
SMPL-X Human Body Mesh Reconstruction. Similar to MANO in hand pose representation, for human body pose representation, we also have SMPL-X [57] representation. It is composed of a set of parameters $(\theta,R,l,\theta_{\text{left}},\theta_{\text{right}})$, where the body pose $\theta\in\mathbb{R}^{N\times 21\times 3}$ represents the $21$ joint orientations(except root), $R,l\in\mathbb{R}^{N\times 3}$ are global orientation and translation, and $\theta_{\text{left}},\theta_{\text{right}}\in\mathbb{R}^{N\times 12}$ are compressed representations of two hands. Human shape parameters $\beta\in\mathbb{R}^{10}$ are given, while $\hat{x}\in\mathbb{R}^{N\times 22\times 3}$ denotes the joint positions we predict.
Due to the flexibility of human body poses, to reconstruct the human body mesh, we not only need joint positions but also joint orientations. Therefore, we need to introduce additional $N\times 21\times 3$ degrees of freedom to every $x_{h_{i}}(i\in[1,n])$, representing the predicted human body pose $\hat{\theta}$ under SMPL-X. These degrees of freedom are also predicted by the diffusion model, but they only participate in the decomposition process and do not get involved in our two synchronization mechanisms.
Here
$\mathcal{L}_{\text{vel}}=\sum\limits_{t=1}^{T-1}\|l_{t}-l_{t+1}\|_{2}^{2}$
which keeps the trajectory smooth.
The hyperparameters involved in the mesh reconstruction process can be found in Table  10 and  11.
@@ C.2 Post Process: Mesh Reconstruction
Table 10: Hyperparameters for MANO hand mesh reconstruction.
Parameter CORE4D BEHAVE
@@
Epoch $5$k $2$k
$\lambda_{\text{pos}}$ $3$ $1$
$\lambda_{\text{vel}}$ $0.1$ $0.1$
Table 11: Hyperparameters for SMPL-X human body mesh reconstruction.
C.3 A Brief Introduction to BPS Algorithm
C.4 Details for Ablation Studies
In our experiments in Section  4, there are three categories of ablation studies: “w/o decompose”, “w/o $\mathcal{L}_{\text{align}}$”, and “w/o exp sync”.
“w/o decompose” means that we direct concatenate the condition vector with $x_{t}$, which substitute the position of $x_{t,\text{dc}}$ in the pipeline figure, and eliminate the branch of $x_{\text{F}}$ or $x_{\text{ac}}$. Note that $x_{t}\neq x_{t,\text{dc}}+x_{t,\text{ac}}$, as $x_{t}$ includes the components with frequencies higher than $\phi_{L}/2\pi$, where $\phi_{L}=L/N$, and $L$ is the cutoff boundary.
“w/o $\mathcal{L}_{\text{align}}$” means that we eliminate the term $\lambda_{\text{align}}\mathcal{L}_{\text{align}}$ in total loss.
“w/o exp sync” means that we perform normal denoising steps every time, without explicit synchronization steps. This can be equivalently interpreted as $s=+\infty$.
C.5 Pseudocode for CSR, CRR, and CSIoU
In this section, we will provide a detailed description for three contact-based metrics: CSR, CSIoU, and CRR.
First, we define two types of contact: surface contact and root contact. Here $o$ is one single object mesh sequence of size $\mathbb{R}^{N\times M\times 3}$, where $M$ is the number of vertices on its mesh. $h$ is hand mesh sequence of size $\mathbb{R}^{N\times 778\times 3}$ in Contact_Surface, while it denotes trajectories of root joints of two hands of size $R^{N\times 2\times 3}$ in Contact_Root, as shown in Algorithm 2.
@@ C.5 Pseudocode for CSR, CRR, and CSIoU
Appendix D Important Statistics
D.1 Dataset Statistics
The sizes of each dataset split are shown in Table  12.
Dataset Statistics
TACO train:test1:test2:test3:test4
=1035:238:260:403:610
CORE4D train:test1:test2=483:197:195
OAKINK2 train:val:test=1884:167:723
GRAB train:val:test=992:198:144 (Unseen Subject)
train:test=1126:208 (Unseen Object)
BEHAVE train:test=231:90
Table 12: Dataset statistics.
D.2 Hyperparameters in Model Architecture
Action/object Label Feature Extraction Branch. Action/object label features are first encoded by pretrained CLIP [64], and then pass through a 2-layer MLP. Specific parameters are shown in Table  13.
Component Description
CLIP Type ViT-B/32
Raw Feature Space Dimension $512$
MLP Architecture Linear($512$, $512$)
ReLU()
Linear($512$, $128$)
Table 13: Hyperparameters in label feature extraction branch.
Object Geometry Feature Extraction Branch. Object geometry features are first encoded by pretrained BPS [59], and then pass through a 2-layer MLP. Specific parameters are shown in Table  14.
Component Description
Raw Feature Space Dimension $1024\times 3$
MLP Architecture Linear($1024\times 3$, $512$)
ReLU()
Linear($512$, $128$)
Table 14: Hyperparameters in rigid body geometry feature extraction branch.
Noise Timestep Embedding Module. Detailed architecture is shown in Table  15.
Component Description
Timestep Embedder SinusoidalPosEmbedding($64$)
Linear($64$, $256$)
GeLU()
Linear($256$, $1024$)
Table 15: Hyperparameters in noise timestep embedding.
Transformer Encoder-Decoder. Concatenate action label features, object label features (for $m$ rigid objects), object geometry features (for $m$ rigid objects), and the shape parameters $\beta$ of $n$ articulated skeletons to form a condition vector, whose shape is $\mathbb{R}^{128\times(2m+1)+10n}$, as is mentioned in Section  3.4. After replicating and concatenating with padding mask, $x_{\text{dc}}$ and $x_{\text{F}}$, the dimension becomes $\mathbb{R}^{N\times\left(128\texttimes(2m+1)+10n+7m+3Dn+7m(m-1)+3Dmn+1\right)}=\mathbb{R}^{N\times\left(128\texttimes(2m+1)+10n+D_{\text{sum}}+1\right)}=\mathbb{R}^{N\times\mathcal{C}}$. The architecture of the transformer encoder-decoder is shown in Table  16.
Component Description
Encoder Conv1D($\mathcal{C}$, $512$)
Latent Transformer $4$-layer, $8$-head, $1024$-dim
Decoder Linear($512$, $D_{\text{sum}}$)
Table 16: Hyperparameters in the transformer encoder-decoder.
Note that after $x_{\text{dc}}$ and $x_{\text{F}}$ (concatenating with the condition vector) passes through the encoder, their shapes becomes $N\times 512$. Concatenating them together results in a latent vector of shape $N\times 1024$, which can be affiliated by the noise embedding, whose dimension is $1024$.
D.3 Training and Inference Hyperparameters
Other important hyperparameters for training and inference process are shown in Table  17.
Parameter TACO CORE4D OAKINK2 GRAB BEHAVE
Batch Size $32$
[…]
Epoch $250$k $140$k $280$k $100$k $100$k
$\lambda_{\text{DC}}$ $1$ $1$ $1$ $1$ $1$
$\lambda_{\text{AC}}$ $2.5$ $0.8$ $0.8$ $0.3$ $0.8$
$\lambda_{\text{norm}}$ $0.1$ $0.1$ $0.1$ $0.1$ $0.1$
$\lambda_{\text{align}}$ $0.3$ $0.3$ $0.3$ $0.15$ $0.15$
$\lambda_{\text{exp}}$ $0.3$ $0.3$ $0.3$ $0.3$ $0.3$
Diffusion $\alpha\in[0.0001,0.01]$, Uniform
D.4 Time and Space Cost, Hardware Configurations
Time and Space Cost for Training and Inference. The training time, average inference time per sample, and GPU memory usage during training are detailed in Table  18.
Cost TACO CORE4D OAKINK2 GRAB BEHAVE
Training $20.7$h $9.5$h $40.1$h $7.9$h $3.7$h
Inference $7.7$s $6.5$s $7.3$s $6.6$s $4.2$s
Memory $6.11$G $5.07$G $8.59$G $4.93$G $3.67$G
Time and Space Cost for Mesh Reconstruction. Time and space cost of performing mesh reconstruction for a motion sequence of $N=200$ frames are shown in Table  19.
Cost TACO CORE4D OAKINK2 GRAB BEHAVE
Time $130$s $144$s $206$s $130$s $74$s
Memory $492$M $826$M $492$M $492$M $430$M
@@ E.1 Current Limitations and Potential Solutions
Some current limitations of SyncDiff and their potential solutions are as follows:
@@
1. Lack of Articulation-Aware Modeling. Our method models articulated objects (such as those in OAKINK2 [108]) as part-wise rigid body individuals directly and coordinates their motions without leveraging their intrinsic articulations. Integrating these articulations into multi-body likelihood modeling could be an interesting future direction. The potential solution may be unify them with the articulated skeletons (hands/humans), and treat their intrinsic articulations as conditions.
@@
2. High-cost of Explicit Synchronization Step. As body number increases, the time consumption for the calculation of alignment loss and explicit synchronization step grows quadratically. Note that for multi-body HOI synthesis, not all pairwise relationships are necessary. A possible solution is to use human priors or another algorithm to filter out the relationships that truly require synchronization, and perform synchronization only across them.
@@
3. Lack of Physically Accurate Guarantees. Unlike methods that utilize true physical simulations, our approach cannot guarantee physical truthfulness. In many cases, minor errors can be observed in the supplementary videos, but these small discrepancies may be sufficient to cause visible failures in real tasks. In robot manipulation tasks, we prefer to treat SyncDiff as a robot planning method that requires downstream integration with physically accurate optimization (RL/IL) to ensure practical usability in real-world applications.
4. Limitations in Pure Multi-human Interaction Synthesis. Since the relative representations need to be generated in the coordinate systems of rigid bodies, whose motions can be represented by translations and rotations, our method may not be directly adapted for pure multi-human interaction synthesis. To address this limitation, more complex relative representations is required. A potential solution is use representations similar to MANO [68], representing the motions of an articulated skeleton as the transformation matrix and local relative positions of its joints. Although this approach sacrifices the homogeneity of the articulated skeleton motion representation, it allows for a more convenient computation of the relative representation between two skeletons. This idea is manifested in a bunch of multi-human interaction synthesis works like ComMDM [71] and InterGen [44].
E.2 Discussions and Some Extensions
This section aims to enumerate some points where readers may have questions or misconceptions, providing detailed explanations. For scenarios where SyncDiff can be extended, we will also offer intuitive methods for expansion.
1. Does SyncDiff lose the flexibility of “one model for all datasets”?
Due to the scarcity of mocap datasets, as well as the need to adopt models into multiple scenarios, sometimes it is necessary to merge data from multiple datasets for training. However, due to the fixed number of parameters, a trained model of a specific size can only handle a pair of (number of rigid objects, number of articulated skeletons). For instance, the data from TACO can only be merged with the 2-hand 2-object samples in OAKINK2 for training, but not directly with GRAB.
Our response is divided into three aspects.
First, the extension from a unified framework to “one model for all datasets” is basically trivial, only increasing some computational load. It is noted that any graphical model is a subgraph of a sufficiently large complete graph. By using masks to cover the individual bodies that do not need to be synthesized and the relative relations that do not need to be adopted, this can be achieved. Such an idea is also reflected in  [72].
Second, existing sota methods can only train for specific (number of rigid objects, number of articulated skeletons), and even require the type of articulated skeleton to be given (for example, the carefully designed grasp guidance in DiffH2O is almost ineffective when applied to the human-object interaction dataset CORE4D; the cross-attention between bodies and contact maps in CG-HOI is too coarse-grained for dexterous hand-object interaction). Our method, as a unified framework, significantly reduces the cost of manually trying different pipelines and designing different guidance/representations for different settings, which is a giant leap forward.
Third, even if our method is treated as a single-track method for every specific configuration, it surpasses many sota methods in the corresponding tracks. As body number increases, motion semantics become more complex, especially between multiple rigid objects. For example, the object trajectories in the GRAB/BEHAVE dataset mainly consist of basic units like picking up, putting down, and lateral movements. However, in TACO, there may be actions such as rubbing back and forth, tapping, and pouring between two items. If OAKINK2 involves modeling the rigid parts of articulated objects, it also needs to ensure coordination between these different parts, such as a test tube passing precisely between the two halves of a test tube holder. As the first work to adopt synchronization and frequency decomposition to model complex motions, we have also significantly outperformed existing state-of-the-art methods in settings with a larger number of bodies.
2. Why doesn’t SyncDiff use the 6 DoF rotation representation that is widely employed in the current motion synthesis methods?
To represent the rotation $R\in SO(3)$ of an object, the 6 DoF representation proposed in  [126] concatenates the first two columns of the $3\times 3$ rotation matrix into a 6-dimensional vector $r$ as the representation. After the model generates the predicted result $\hat{r}$, it is split into two column vectors, normalized, and subjected to Gram-Schmidt orthogonalization, and then completed into the rotation matrix $\hat{R}$ in $SO(3)$.
In our method, since the relative representations need to be solved frequently, if we use such 6 DoF representation, it is necessary to complete them into matrices for operations such as inversion and composition, which significantly increases the computational cost. Therefore, we adopt quaternions, which are more convenient for normalization and calculation.
3. Is the explicit synchronization during inference in SyncDiff inspired by the guidance strategy in Guided Motion Diffusion (GMD) [37]?
As is stated in Section  2.2, several prominent works inject external priors or constraints into synthesized results by performing linear fusions, imputations or inpaintings during the inference process of diffusion models. The term $-\nabla_{\textbf{x}_{0}}G_{z}(P_{x}^{z}\textbf{x})$ in GMD and the gradient of negative log-likelihood (actually equivalent to the score functions) in SyncDiff (Refer to Section  A.2) both provide the direction of denoising (for both), adhering to constraints (for GMD), or synchronization (for SyncDiff).
The distinction lies in that different application scenarios have led to different concrete operations for the same idea. GMD aims to constrain the generation of high-dimensional human pose $x$ controlled by the goal function $G_{z}$, where $z$ is low-dimensional pelvis trajectory. Hence, it employs an inpainting/imputation strategy to project (by the projector $P_{x}^{z}$) and complete the sparse trajectory during inference. In contrast, in SyncDiff, “synchronization” is a concept that is difficult to quantify. Therefore, we introduce the probabilistic modeling on graphical models, and ensure that synchronization and data denoising are simultaneously established in the inference process.
4. Can the articulated skeletons in SyncDiff be other body parts besides hands and the full body?
In Section  3.1, our definition of articulated skeletons is “a set of joints whose motion can be reconstructed based on joint information and shape parameter $\beta$”. To extend SyncDiff to HOI motion synthesis of other body parts, two issues need to be addressed:
1) Collect relevant high-precision mocap data. For example, if we want to perform HOI motion synthesis for feet, we need to collect high-quality data such as kicking a ball, putting on shoes, ice skating, etc.
2) Design a data representation for motion reconstruction based on joint information and shape parameter $\beta$, similar to MANO for hands or SMPL-X for human bodies.
5. Does SyncDiff’s ignorance of affordance result in the item being manipulated incorrectly? For example, the cup was not picked up from the handle position?
SyncDiff does not consider affordance because the five datasets we use do not provide affordance for rigid objects at all. If the affordance needs to be considered, the relevant data with affordance map needs to be collected first, and the next potential solution may be to encode it in some way as part of the object’s geometric features.
In addition, the data-driven generative model should fit the data distribution. As long as the dataset used for training ensures high-quality and correct interactions in a wide range, the model should learn to generate similar samples with correct manipulation by itself.